# Commutativity criteria for prime rings with involution via pairs of endomorphisms

### v1

@@ 2. Main Results
[11, Fact 1] Let $R$ be a 2-torsion free non-commutative prime ring with involution of the second kind.
  • (i)
    If $a[x,x^{\ast}]b=0$ for all $x\in R$; then $a=0$ or $b=0$.
  • (ii)
    If $a(x\circ x^{\ast})b=0$ for all $x\in R$; then $a=0$ or $b=0$.
Fact 2.4.
@@
Throughout this paper, by a nontrivial endomorphism, we mean an endomorphism that is not the identity mapping on the given ring.
@@
For some $h\in H(R)\cap Z(R),$ taking $x=h$ in the above relation, we obtain $g_{1}(h)g_{2}(h)-h^{2}\in Z(R).$ Thus, we also get $g_{1}(h)g_{2}(h)\in Z(R).$ Replacing $x$ by $xh$ in (2.1) in order to see that
@@
Using primeness of $R$ we obtain that either $g_{1}(x)g_{2}(x^{\ast})\in Z(R)$ for all $x\in R$ or $g_{1}(h)g_{2}(h)=h^{2}$ for all $h\in H(R)\cap Z(R).$ In the first case it follows from our hypothesis that $xx^{\ast}\in Z(R)$ for all $x\in R.$ Replacing $x$ by $x^{\ast}$ in the last relation, we find $x^{\ast}x\in Z(R)$ for all $x\in R.$ It forces that $x,x^{\ast}]\in Z(R)$ for all $x\in R,$ hence $R$ is commutative by Fact 2.4.
@@
Replacing $x$ by $xh$ in the last obtained expression, we may infer than
@@
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{1})$ with $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R),$ then $R$ is commutative.
@@
Let us suppose that
@@
That means $g_{1}(h)=g_{2}(h)=h$ for all $h\in H(R)\cap Z(R).$ Replacing $h$ by $k_{c}^{2},$ where $k_{c}\in S(R)\cap Z(R),$ we get $g_{1}(k_{c})^{2}=k_{c}^{2},$ which implies that either $g_{1}(k_{c})=k_{c}$ or $g_{1}(k_{c})=-k_{c}.$ Similarly for all $k_{c}\in S(R)\cap Z(R),$ we get $g_{2}(k_{c})=k_{c}$ or $g_{2}(k_{c})=-k_{c}.$ Thus, we have the following cases:
@@
Let us discuss these cases in detail and see their respective outcomes. First let us polarizing our hypothesis to get
@@
(ii) Suppose that $g_{1}(k_{c})=-k_{c}$ and $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ It is not difficult to see that this case also leads us to our conclusion by following the same reasoning as of case (i).
@@
Finally, it can be seen that (iv) follows in the same manner as (iii), so we omit its proof. ∎
@@
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{1}),$ then $R$ is a commutative.
@@
Suppose that $R$ is not commutative. Then by Lemma 2.5, we know that $g_{1}$ and $g_{2}$ both are center-preserving. We shall now use this fact implicitly in order to proceed further. Replacing $x$ by $xh$ in (2.3), we have
@@
Case (c). This case is done by Lemma 2.6. Thus we conclude that $R$ is commutative, as desired. ∎
@@
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{2})$ with $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R),$ then $R$ is commutative.
@@
The proof proceeds almost in the same manner as that of Lemma 2.6, except for some variations in the mapping $g_{2}$, which we shall now examine. Following in the same way we get that $g_{2}(h)=-h$ for all $h\in H(R)\cap Z(R)$. Moreover, for each $k_{c}\in S(R)\cap Z(R),$ either $g_{1}(k_{c})=k_{c}$ or $g_{1}(k_{c})=-k_{c}.$ But clearly, can not explicitly have the same behavior of $g_{2}$ on central skew-symmetric elements. In order to obtain that, we need more reasoning. Polarizing the given hypothesis, we have
@@
In case $g_{2}(R)\subseteq Z(R)$ from (2.25), we observe that $g_{1}(x)+x+x^{\ast}\in Z(R)$ for all $x\in R.$ For some $0\neq k_{c},$ putting $x=xk_{c}$ in it, we get $g_{1}(x)+x-x^{\ast}\in Z(R)$ for all $x\in R.$ Comparing last two expressions obtained, we find that $2x^{\ast}\in Z(R)$ for all $x\in R,$ it forces $R$ to be commutative, and hence we are done. On the other hand, we have the following two cases:
Case 1. Let $g_{2}(k_{c})=k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ Since it is well known that for any $h\in H(R)\cap Z(R),$ $hk_{c}\in S(R)\cap Z(R).$ In view of our assumption, we see that
$hk_{c}$ $=g_{2}(hk_{c})$
$=g_{2}(h)g_{2}(k_{c})$
$=-hk_{c}.$
It forces $hk_{c}=0,$ implying $h=0$ for all $h\in H(R)\cap Z(R)$ or $k_{c}=0$ for all $k_{c}\in S(R)\cap Z(R).$ Thus it leads us to a contradiction.
Case 2. Let $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ In this case, we get contradiction in the same way. ∎
@@
The proof follows on the same lines of Theorem 2.7
With some insignificant variation, one can easily observe the following theorems.
Theorem 2.10.
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{3}),$ then $R$ is a commutative.
Proof.
The proof proceeds along the same lines as that of Theorem 2.7. ∎
Theorem 2.11.
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{4}),$ then $R$ is a commutative.
Proof.
The proof proceeds along the same lines as that of Theorem 2.9. ∎
The following example exhibits that the condition on involution $\ast$ is of the second kind in Theorem 2.7 and 2.9 is not superfluous.
@@
In 2020, Mamouni et al. [11] showed that if a 2-torsion free prime ring $R$ with involution $\ast$ of the second kind, admits a nontrivial endomorphism $g$ satisfying
$g(xx^{\ast})-g(x^{\ast}x)=[x,x^{\ast}]\quad\text{or}\quad g(xx^{\ast})+g(x^{\ast}x)=x\circ x^{\ast}$
for all $x\in R,$ then $R$ must be commutative. It is natural and of considerable interest to conjecture that such a ring is commutative in the presence of two distinct endomorphisms involved in the aforementioned identities. In this section, we address this conjecture and establish its validity by providing a positive solution.
Theorem 2.14.
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ such that
$g_{1}(xx^{\ast})-g_{2}(x^{\ast}x)=[x,x^{\ast}],~\forall~x\in R,$
then $R$ is a commutative.
Proof.
Let us suppose that
(2.31) $g_{1}(xx^{\ast})-g_{2}(x^{\ast}x)=[x,x^{\ast}],~\forall~x\in R.$
Linearizing this relation, we see that
(2.32) $g_{1}(xy^{\ast})+g_{1}(yx^{\ast})-g_{2}(x^{\ast}y)-g_{2}(y^{\ast}x)=[x,y^{\ast}]+[y,x^{\ast}],~\forall~x,y\in R.$
Replacing $y$ by $yh$ in (2.32), where $h\in H(R)\cap Z(R),$ we get
(2.33) $g_{1}(xy^{\ast})g_{1}(h)+g_{1}(yx^{\ast})g_{1}(h)-g_{2}(x^{\ast}y)g_{2}(h)-g_{2}(y^{\ast}x)g_{2}(h)=([x,y^{\ast}]+[y,x^{\ast}])h.$
Right multiplying (2.32) by $g_{1}(h)$ and then subtracting it from (2.33) in order to find that
(2.34) $(g_{2}(x^{\ast}y)+g_{2}(y^{\ast}x))(g_{2}(h)-g_{1}(h))=([x,y^{\ast}]+[y,x^{\ast}])(h-g_{1}(h)).$
Taking $yh$ for $y$ in (2.34), we have
(2.35) $g_{2}(h)(g_{2}(x^{\ast}y)+g_{2}(y^{\ast}x))(g_{2}(h)-g_{1}(h))=h([x,y^{\ast}]+[y,x^{\ast}])(h-g_{1}(h)).$
Left multiplying (2.34) with $g_{2}(h)$ and then subtracting it from (2.35), we obtain $(g_{2}(h)-h)([x,y^{\ast}]+[y,x^{\ast}])(h-g_{1}(h))=0$ for all $x,y\in R$ and $h\in H(R)\cap Z(R).$ In particular, we get $(g_{2}(h)-h)([x,x^{\ast}](g_{1}(h)-h)=0$ for all $x,y\in R$ and $h\in H(R)\cap Z(R).$ Invoking Fact 2.3, we get either $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R)$ or $g_{2}(h)=h$ for all $h\in H(R)\cap Z(R)$.
Let us first consider $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R).$ Then in view of (2.34), we find $(g_{2}(x^{\ast}y)+g_{2}(y^{\ast}x))(g_{2}(h)-h)=0$ for all $x,y\in R$ and $h\in H(R)\cap Z(R).$ In particular, we find $g_{2}(x^{\ast}x)(g_{2}(h)-h)=0$ for all $x,y\in R$ and $h\in H(R)\cap Z(R).$ Now it is immediate from (2.31) that $g_{1}(xx^{\ast})(g_{2}(h)-h)=[x,x^{\ast}](g_{2}(h)-h)$ for all $x\in R.$ Putting $x=h^{\prime},$ where $0\neq h^{\prime}\in H(R)\cap Z(R),$ we get $g_{1}(h^{\prime 2})(g_{2}(h)-h)=0$ for all $h\in H(R)\cap Z(R).$ This implies that $h^{\prime 2}(g_{2}(h)-h)=0$ for all $h\in H(R)\cap Z(R).$ It forces that $g_{2}(h)=h$ for all $h\in H(R)\cap Z(R).$
On the other hand, we are assuming that $g_{2}(h)=h$ for all $h\in H(R)\cap Z(R).$ Thus from (2.34), we see that
$(g_{2}(x^{\ast}y)+g_{2}(y^{\ast}x))(h-g_{1}(h))=([x,y^{\ast}]+[y,x^{\ast}])(h-g_{1}(h)).$
For $0\neq h^{\prime}\in H(R)\cap Z(R),$ substituting $x=h^{\prime}=y$ in the above relation, we get $h^{\prime 2}(h-g_{1}(h))=0$ for all $h\in H(R)\cap Z(R),$ which yields that $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R).$
Thus with this, we conclude that $g_{1}(h)=h$ if and only if $g_{2}(h)=h$ for any $h\in H(R)\cap Z(R).$ Now, let us substitute $y=h$ in (2.32), we find that
(2.36) $g_{1}(x+x^{\ast})-g_{2}(x+x^{\ast})=0,~\forall~x\in R.$
Putting $x=xy^{\ast}$ in (2.36), we see that
(2.37) $g_{1}(xy^{\ast}+yx^{\ast})=g_{2}(xy^{\ast}+yx^{\ast}),~\forall~x,y\in R.$
Using (2.37) in (2.32), we find
$g_{2}(xy^{\ast})+g_{2}(yx^{\ast})-g_{2}(x^{\ast}y)-g_{2}(y^{\ast}x)=[x,y^{\ast}]+[y,x^{\ast}],~\forall~x,y\in R.$
It implies that
$g_{2}([x,y^{\ast}])+g_{2}([y,x^{\ast}])=[x,y^{\ast}]+[y,x^{\ast}],~\forall~x,y\in R.$
In particular, it gives us the relation $g_{2}([x,x^{\ast}])=[x,x^{\ast}]$ for all $x\in R.$ Thus the conclusion follows from Theorem 1 of [11]. ∎
Example 2.15.
Let $S$ be a commutative domain and $\xi:S\to S$ be any endomorphism. Let $P$ be a noncommutative ring with involution #\# of the second kind. Let $R=S\times P$ and $\ast$ be the involution of $R$ defined as $\ast=(id_{S},\#),$ which is of the second kind. Define an endomorphism $g:R\to R$ such that $g(x,y)=(\xi(x),y).$ Then it can be seen that for $g_{1}=g$ and $g_{2}=g$ the identity $g_{1}(XX^{\ast})-g_{2}(X^{\ast}X)=[X,X^{\ast}],$ but $R$ is not commutative.
By taking $g_{1}=g_{2}$ in Theorem 2.14, we get the following result:
Corollary 2.16.
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphism $g$ such that
$g(x),g(x^{\ast})]=[x,x^{\ast}],~\forall~x\in R,$
then $R$ is a commutative.
Theorem 2.17.
[…]
$g_{1}(xx^{\ast})+g_{2}(x^{\ast}x)=x\circ x^{\ast},~\forall~x\in R,$
then $R$ is a commutative.
Proof.
Let us suppose that
(2.38) $g_{1}(xx^{\ast})+g_{2}(x^{\ast}x)=x\circ x^{\ast},~\forall~x\in R.$
Linearizing this expression, we find that
(2.39) $g_{1}(xy^{\ast})+g_{1}(yx^{\ast})+g_{2}(x^{\ast}y)+g_{2}(y^{\ast}x)=x\circ y^{\ast}+y\circ x^{\ast},~\forall~x,y\in R.$
Replacing $y$ by $yh$ in (2.39), where $h\in H(R)\cap Z(R)$, we have
(2.40) $g_{1}(x^{\ast}y+y^{\ast}x)g_{1}(h)+g_{2}(xy^{\ast}+yx^{\ast})g_{2}(h)=(x\circ y^{\ast}+y\circ x^{\ast})h,~\forall~x,y\in R.$
Right multiplying (2.39) by $g_{1}(h)$ and then subtracting it from (2.40), we obtain
(2.41) $g_{2}(xy^{\ast}+yx^{\ast})(g_{2}(h)-g_{1}(h))=(x\circ y^{\ast}+y\circ x^{\ast})(h-g_{1}(h).$
Replacing $y$ by $yh$ in (2.41), where $h\in H(R)\cap Z(R)$ and then using it, we conclude that
(2.42) $g_{2}(h)g_{2}(xy^{\ast}+yx^{\ast})(g_{2}(h)-g_{1}(h))=h(x\circ y^{\ast}+y\circ x^{\ast})(h-g_{1}(h).$
Comparing (2.41) in (2.42), we get
(2.43) $(g_{2}(h)-h)(x\circ y^{\ast}+y\circ x^{\ast})(h-g_{1}(h)=0\mbox{~~for~all~~}x,y\in R,\;h\in H(R)\cap Z(R).$
It implies that $(g_{2}(h)-h)(x\circ x^{\ast})(g_{1}(h)-h)=0$ for all $x,y\in R,\;h\in H(R)\cap Z(R)$. Invoking Fact 2.3, we have either $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R)$ or $g_{2}(h)=h$ for all $h\in H(R)\cap Z(R)$.
Now, if $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R)$, then (2.38) implies $g_{2}(h)^{2}=h^{2}$ for all $h\in H(R)\cap Z(R)$. For some $0\neq h^{\prime}\in H(R)\cap Z(R),$ putting $x=y=h^{\prime}$ in (2.41) and then using $g_{1}(h)=h$ and $g_{2}(h)^{2}=h^{2}$, we find $g_{2}(h)=h$ for all $h\in H(R)\cap Z(R).$
In the same way, if $g_{2}(h)=h$ for all $h\in H(R)\cap Z(R)$, then taking $x=y=h^{\prime}$ in (2.41), we get $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R).$ Thus we conclude that $g_{2}(h)=h$ if and only if $g_{1}(h)=h$ for any $h\in H(R)\cap Z(R)$.
Taking $y=h,$ where $0\neq h\in H(R)\cap Z(R)$ in (2.39), we have
(2.44) $g_{1}(x+x^{\ast})+g_{2}(x+x^{\ast})=2(x+x^{\ast}),~\forall~x\in R.$
Substituting $x=xy^{\ast}$ in (2.44), we get
(2.45) $g_{1}(xy^{\ast}+yx^{\ast})+g_{2}(xy^{\ast}+yx^{\ast})=2(xy^{\ast}+yx^{\ast}),~\forall~x,y\in R.$
Subtracting (2.39) from (2.45), we have
$g_{2}([x,y^{\ast}])+g_{2}([y,x^{\ast}])=[x,y^{\ast}]+[y,x^{\ast}],~\forall~x,y\in R.$
In particular, we obtain $g_{2}(x),g_{2}(x^{\ast})]=[x,x^{\ast}]$ for all $x\in R.$ Again by Theorem 1 of [11], $R$ is commutative. ∎
By taking $g_{1}=g_{2}$ in Theorem 2.17, we get the following result:
Corollary 2.18.
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphism $g$ such that
$g(x)\circ g(x^{\ast})=x\circ x^{\ast},~\forall~x\in R,$
then $R$ is a commutative.
We shall end our discussion with the following examples.
Example 2.19.
Let us consider $R=M_{2}(\mathbb{C})\times\mathbb{C}$ be a nocommutative ring which is not prime. Let us define an involution #\# on $M_{2}(\mathbb{C})$ as $\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)^{\#}=\left(\begin{array}[]{cc}\overline{d}&\overline{b}\\ \overline{c}&\overline{a}\end{array}\right),$ which is the second kind and an endomorphism $f\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)=\left(\begin{array}[]{cc}a&-b\\ -c&d\end{array}\right)$ on $R.$ Thus we can see that a function $g:R\to R$ such that $g(X,u)=(f(X),id_{\mathbb{C}}(u))$ is a non-identity endomorphism of $R$ and $(X,u)^{\ast}=(X^{\#},\overline{u})$ is an involution of the second kind on $R.$ Now, it can be easily observed that if $g_{1}=g$ and $g_{2}=g,$ then $g_{1}(xx^{\ast})+g_{2}(x^{\ast}x)=x\circ x^{\ast}$ for all $x\in R,$ but $R$ is not commutative. Therefore, the assumption of $R$ a prime ring is not redundant in Theorem 2.17.
Example 2.20.
Let $R=\left\{\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)|~a,b,c,d\in\mathbb{Q}\right\}$, where $\mathbb{Q}$ denotes the ring of rational numbers with involution as in Example 2.12. Define the mappings $g_{1},g_{2}:R\to R$ by $g_{1}\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)=\left(\begin{array}[]{cc}d&c\\ b&a\end{array}\right),$ $g_{2}\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)=\left(\begin{array}[]{cc}a&-b\\ -c&d\end{array}\right)$ It is straight forward to see that $R$ is prime ring with involution $\ast$ and non-identity endomorphisms $g_{1}$ and $g_{2}.$ Clearly, $\ast$ is involution of the second kind and the following assertions hold:
  • (i)
    $g_{1}(xx^{\ast})-g_{2}(x^{\ast}x)=[x,x^{\ast}]$ for all $x\in R;$
  • (ii)
    $g_{1}(xx^{\ast})+g_{2}(x^{\ast}x)=x\circ x^{\ast}$ for all $x\in R.$
But, $R$ is not commutative. Thus, the assumption of $\ast$ being the second kind involution is essential in Theorem 2.14 and 2.17.

### v2

@@ 2. Main Results
Throughout this paper, by a nontrivial endomorphism, we mean an endomorphism which is not identity on the given ring.
@@
Linearizing this relation, we get
For some $h\in H(R)\cap Z(R),$ taking $x=h$ in the relation (2.1), we obtain $g_{1}(h)g_{2}(h)-h^{2}\in Z(R),$ which forces $g_{1}(h)g_{2}(h)\in Z(R).$ Replacing $x$ by $xh$ in (2.1), we have
@@
Using primeness of $R$ we obtain that either $g_{1}(x)g_{2}(x^{\ast})\in Z(R)$ for all $x\in R$ or $g_{1}(h)g_{2}(h)=h^{2}$ for all $h\in H(R)\cap Z(R).$ In the first case it follows from our hypothesis that $xx^{\ast}\in Z(R)$ for all $x\in R.$ Replacing $x$ by $x^{\ast}$ in the last relation, we find $x^{\ast}x\in Z(R)$ for all $x\in R.$ It forces that $x,x^{\ast}]\in Z(R)$ for all $x\in R,$ hence $R$ is commutative by Fact 2.3.
@@
Replacing $x$ by $xh$ in the last obtained expression, we may infer that
@@
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{1})$ or $(A_{3})$ with $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R),$ then $R$ is commutative.
@@
Let assume that $(A_{1})$ holds. Suppose $R$ is not commutative. By hypothesis, we have
@@
This means that $g_{1}(h)=g_{2}(h)=h$ for all $h\in H(R)\cap Z(R).$ Replacing $h$ by $k_{c}^{2},$ where $k_{c}\in S(R)\cap Z(R),$ we get $g_{1}(k_{c})^{2}=k_{c}^{2}.$ It takes us to $(g_{1}(k_{c})-k_{c})(g_{1}(k_{c})+k_{c})=0.$ Invoking Lemma 2.4, we obtain that either $g_{1}(k_{c})=k_{c}$ or $g_{1}(k_{c})=-k_{c}.$
Similarly for all $k_{c}\in S(R)\cap Z(R),$ we get $g_{2}(k_{c})=k_{c}$ or $g_{2}(k_{c})=-k_{c}.$ Thus, the following cases arise:
@@
Let us discuss these cases in detail and see their respective outcomes.
@@
(ii) Assume that $g_{1}(k_{c})=-k_{c}$ and $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ Substituting $yk_{c}$ for $y$ in (2.12), where $0\neq k_{c}\in S(R)\cap Z(R)$, we have
$\Big(g_{1}(x)g_{2}(y^{\ast})-g_{1}(y)g_{2}(x^{\ast})+xy^{\ast}-yx^{\ast}\Big)k_{c}\in Z(R),~\forall~x,y\in R.$
Since $k_{c}$ is nonzero, by primeness of $R,$ we find that
(2.17) $g_{1}(x)g_{2}(y^{\ast})-g_{1}(y)g_{2}(x^{\ast})+xy^{\ast}-yx^{\ast}\in Z(R),~\forall~x,y\in R.$
Combining (2.26) and (2.34) to conclude that $g_{1}(x)g_{2}(y^{\ast})-yx^{\ast}\in Z(R)$ for all $x,y\in R.$ Taking $x=h,$ where $0\neq h\in H(R)\cap Z(R),$ we get $g_{2}(y^{\ast})-y\in Z(R)$ for all $y\in R.$ If we replace $y$ by $y^{\ast}$ in this relation, then we find an identical expression to Eq. (15) of [12], repeating the same reasoning, we get our conclusion.
@@
(iv) Finally, we consider that $g_{1}(k_{c})=k_{c}$ and $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ It can be seen that this case follows in the same manner as (iii), so we omit its proof.
It completes the proof.
It can be verified that the same conclusion follows from condition $(A_{3})$ by an argument parallel to the above, with only minor modifications. Hence, the proof is omitted. ∎
@@
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{1})$ or $(A_{3}),$ then $R$ is a commutative.
@@
Suppose that $R$ is not commutative. Then by Lemma 2.4, we know that $g_{1}$ and $g_{2}$ both are center-preserving. We shall now use this fact implicitly in order to proceed further and also recall some relations obtain in Lemma 2.4. Replacing $x$ by $xh$ in (2.4), we have
@@
Case (c). This case is done by Lemma 2.5.
Therefore, we conclude that $R$ is commutative, as desired.
It can be verified that the same conclusion follows from condition $(A_{3})$ by an argument parallel to the above, with only minor modifications. Hence, the proof is omitted. ∎
@@
Let $R$ be a 2-torsion free prime ring with involution $\ast$ of the second kind. If $R$ admits nontrivial endomorphisms $g_{1},g_{2}$ satisfying $(A_{2})$ or $(A_{4})$ with $g_{1}(h)=h$ for all $h\in H(R)\cap Z(R),$ then $R$ is commutative.
@@
The proof proceeds almost in the same manner as that of Lemma 2.5, except for some variations in the mapping $g_{2}$, which we shall now examine. Following in the same way we get that $g_{2}(h)=-h$ for all $h\in H(R)\cap Z(R)$. Moreover, for each $k_{c}\in S(R)\cap Z(R),$ either $g_{1}(k_{c})=k_{c}$ or $g_{1}(k_{c})=-k_{c}.$ But we can not explicitly have the same behavior of $g_{2}$ on central skew-symmetric elements. In order to obtain that, we need more reasoning. Polarizing the given hypothesis, we have
@@
First let us discuss the case $g_{2}(R)\subseteq Z(R).$ From (2.27), we observe that $g_{1}(x)+x+x^{\ast}\in Z(R)$ for all $x\in R.$ For some $0\neq k_{c},$ putting $x=xk_{c}$ in it, we get $g_{1}(x)+x-x^{\ast}\in Z(R)$ for all $x\in R.$ Comparing last two expressions obtained, we find that $2x^{\ast}\in Z(R)$ for all $x\in R,$ it forces $R$ to be commutative, and hence we are done. On the other hand, we have the following cases:
  • (i)
    $g_{1}(k_{c})=k_{c}$ and $g_{2}(k_{c})=k_{c}$ for all $k_{c}\in S(R)\cap Z(R);$
  • (ii)
    $g_{1}(k_{c})=-k_{c}$ and $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R);$
  • (iii)
    $g_{1}(k_{c})=-k_{c}$ and $g_{2}(k_{c})=k_{c}$ for all $k_{c}\in S(R)\cap Z(R);$
  • (iv)
    $g_{1}(k_{c})=k_{c}$ and $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$
(i) Let us suppose that $g_{1}(k_{c})=k_{c}$ and $g_{2}(k_{c})=k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ Placing $yk_{c}$ instead of $y$ in (2.26), where $0\neq k_{c}\in S(R)\cap Z(R)$, we have
$\Big(-g_{1}(x)g_{2}(y^{\ast})+g_{1}(y)g_{2}(x^{\ast})-xy^{\ast}+yx^{\ast}\Big)k_{c}\in Z(R),~\forall~x,y\in R.$
Since $k_{c}$ is nonzero, by primeness of $R,$ we find that
(2.33) $g_{1}(x)g_{2}(y^{\ast})+g_{1}(y)g_{2}(x^{\ast})-xy^{\ast}+yx^{\ast}\in Z(R),~\forall~x,y\in R.$
Jointly considering (2.26) and (2.33) to conclude that $g_{1}(y)g_{2}(x^{\ast})+yx^{\ast}\in Z(R)$ for all $x,y\in R.$ Replacing $x$ by $x^{\ast},$ we obtain $g_{1}(y)g_{2}(x)+yx\in Z(R)$ for all $x,y\in R.$ Hence, the conclusion is obtained by [6, Theorem 2.4].
(ii) Assume that $g_{1}(k_{c})=-k_{c}$ and $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ Substituting $yk_{c}$ for $y$ in (2.26), where $0\neq k_{c}\in S(R)\cap Z(R)$, we have
$\Big(g_{1}(x)g_{2}(y^{\ast})-g_{1}(y)g_{2}(x^{\ast})-xy^{\ast}+yx^{\ast}\Big)k_{c}\in Z(R),~\forall~x,y\in R.$
Since $k_{c}$ is nonzero, by primeness of $R,$ we find that
(2.34) $g_{1}(x)g_{2}(y^{\ast})-g_{1}(y)g_{2}(x^{\ast})-xy^{\ast}+yx^{\ast}\in Z(R),~\forall~x,y\in R.$
Combining (2.26) and (2.34) to conclude that $g_{1}(x)g_{2}(y^{\ast})+yx^{\ast}\in Z(R)$ for all $x,y\in R.$ Taking $x=h,$ where $0\neq h\in H(R)\cap Z(R),$ we get $g_{2}(y^{\ast})+y\in Z(R)$ for all $y\in R.$ If we replace $y$ by $y^{\ast}$ in this relation, then we find an identical expression to Eq. (19) of [12], following in the same way, we get our conclusion.
(iii) Let us consider $g_{1}(k_{c})=-k_{c}$ and $g_{2}(k_{c})=k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ Substituting $yk_{c}$ for $y$ in (2.26), where $0\neq k_{c}\in S(R)\cap Z(R)$, we find that
(2.35) $g_{1}(x)g_{2}(y^{\ast})-g_{1}(y)g_{2}(x^{\ast})-xy^{\ast}+yx^{\ast}\in Z(R),~\forall~x,y\in R.$
On adding this in (2.26), we get $yx^{\ast}\in Z(R)$ for all $x,y\in R.$ It forces commutativity of $R,$ which is a contradiction. Hence we are done.
(iv) finally, let us assume that $g_{1}(k_{c})=k_{c}$ and $g_{2}(k_{c})=-k_{c}$ for all $k_{c}\in S(R)\cap Z(R).$ This case can be easily seen in the same manner of Case (iii).
Therefore, $R$ is commutative.
It can be verified that the same conclusion follows from condition $(A_{4})$ by an argument parallel to the above, with only minor modifications. Hence, the proof is omitted. ∎
@@
The proof follows on the same lines of Theorem 2.6. ∎
The following example exhibits that the condition on involution $\ast$ is of the second kind in Theorem 2.6 and 2.8 is not superfluous.